# Diff-IP2D: Diffusion-Based Hand-Object Interaction Prediction on Egocentric Videos

### v4

@@ III-A Preliminaries
Task definition. Given the video clip of past egocentric observations =$\mathcal{I}=\{I_{t}\}_{t=-N_{\text{p}}+1}^{0}$, we aim to predict future hand trajectories =$\mathcal{H}=\{H_{t}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{L}}}\}_{t=1}^{N_{\text{f}}}(H_{t}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{L}}}\in\mathbb{R}^{2})$ and potential object contact points $\mathcal{O}=\{O_{n}\}_{n=1}^{N_{\text{o}}}(O_{n}\in\mathbb{R}^{2})$, where $N_{\text{p}}$ and $N_{\text{f}}$ are the numbers of frames in the past and future time horizons respectively, and $N_{\text{o}}$ denotes the number of predicted contact points used to calculate interaction hotspots as object affordances. Following the previous works [10, 12], we predict future positions of the right hand, left hand, and affordance of the next active object on the last observed image of the input videos as a canvas.
@@ III-B Architecture
Predictors. Our proposed predictors consist of Hand Trajectory Head (HTH) and Object Affordance Head (OAH). HTH contains an MLP that receives the future parts of the denoised features $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}[N_{\text{p}}\!+\!1\!:\!N_{\text{p}}\!+\!N_{\text{f}}]$ and $\hat{F}^{\scriptstyle{\text{L}}}_{\text{seq}}[N_{\text{p}}\!+\!1\!:\!N_{\text{p}}\!+\!N_{\text{f}}]$, to generate future hand waypoints $\mathcal{H}$ of two hands. As to OAH, we empirically exploit Conditional Variational Autoencoder (C-VAE) [53] to generate possible contact points $\mathcal{O}$ of the next active object in near future. Taking the right hand as an example, the condition is selected as the time-averaged $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}$ and predicted waypoints $H_{t}^{\scriptstyle{\text{R}}}$. Note that we additionally consider denoised future HOI features $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}[N_{\text{p}}\!+\!1:N_{\text{p}}\!+\!N_{\text{f}}]$ ($t\!>\!0$) besides the features from past observations ($t\!\leq\!0$) for object affordance prediction. This aligns with the intuitive relationship between the contact points and the overall interaction process.
@@ III-C Training
$\textstyle\mathcal{L}_{\text{VLB}}^{\scriptstyle{\text{R}}}=\sum_{s=2}^{S}||\mathbf{z}_{0}^{\scriptstyle{\text{R}}}-f_{\scriptscriptstyle\text{MADT}}(\mathbf{z}_{s}^{\scriptstyle{\text{R}}},s,M_{\text{seq}})||^{2}$
$+||F_{\text{seq}}^{\scriptstyle{\text{R}}}-\hat{F}_{\text{seq}}^{\scriptstyle{\text{R}}}||^{2},$ (1)
@@
$\textstyle\mathcal{L}_{\text{traj}}^{\scriptstyle{\text{R}}}=\sum_{t=1}^{N_{\text{f}}}||H_{t}^{\scriptstyle{\text{R}}}-H_{t}^{{\scriptstyle{\text{R}}},\text{gt}}||^{2},$ (2)
@@
$\textstyle\mathcal{L}_{\text{aff}}=\sum_{n=1}^{N_{\text{o}}}||O_{n}-O_{n}^{\text{gt}}||^{2}+c\mathcal{L}_{\text{KL}},$ (3)
@@
$\textstyle\mathcal{L}_{\text{reg}}^{\scriptstyle{\text{R}}}=\sum_{t=1}^{N_{\text{f}}}||\tilde{H}_{t}^{\scriptstyle{\text{R}}}-H_{t}^{{\scriptstyle{\text{R}}},\text{gt}}||^{2},$ (4)
@@ III-D Inference
${}^{*}\!\uparrow$ ${}^{*\!}\uparrow$ ${}^{*\!}\uparrow$ ${}^{*\!}\uparrow$ ${}^{*\!}\uparrow$ ${}^{*\!}\uparrow$ ${}^{*\!}\uparrow$ ${}^{*\!}\uparrow$ ${}^{*\!}\uparrow$
@@ IV-B Separate Evaluation on Hand Trajectory and Object Affordance Prediction
WDE $\downarrow$ FDE $\downarrow$ SIM $\uparrow$ AUC-J $\uparrow$ NSS $\uparrow$ ${}^{*\!}\uparrow$ ${}^{*\!}\uparrow$ ${}^{*\!}\uparrow$
@@ 1 Motion-Related Gaps and Egomotion Homography
      ${}^{*\!}\uparrow$       ${}^{*\!}\uparrow$       ${}^{*\!}\uparrow$       ${}^{*\!}\uparrow$       ${}^{*\!}\uparrow$       ${}^{*\!}\uparrow$
@@ 4 More Details about Datasets and Diff-IP2D Training Configurations
For training Diff-IP2D, we use AdamW optimizer [63] with a learning rate $2$e-$4$. The total loss function is depicted below. The loss weights are initially set as $\lambda_{\text{VLB}}=1$, $\lambda_{\text{traj}}=1$, $\lambda_{\text{aff}}=0.1$, and $\lambda_{\text{reg}}=0.2$. All the networks in Diff-IP2D are trained for $30$ epochs with a batch size of $8$ on 2 A100 GPUs.

### v5

@@ III-A Preliminaries
Task definition. Given the video clip of past egocentric observations $\mathcal{I}=\{I_{t}\}_{t=-N_{\text{p}}+1}^{0}$, we aim to predict future hand trajectories $\mathcal{H}=\{H_{t}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{L}}}\}_{t=1}^{N_{\text{f}}}(H_{t}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{L}}}\in\mathbb{R}^{2})$ and potential object contact points $\mathcal{O}=\{O_{n}\}_{n=1}^{N_{\text{o}}}(O_{n}\in\mathbb{R}^{2})$, where $N_{\text{p}}$ and $N_{\text{f}}$ are the numbers of frames in the past and future time horizons respectively, and $N_{\text{o}}$ denotes the number of predicted contact points used to calculate interaction hotspots as object affordances. Following the previous works [10, 12], we predict future positions of the right hand, left hand, and affordance of the next active object on the last observed image of the input videos as a canvas.
@@ III-B Architecture
Predictors. Our proposed predictors consist of Hand Trajectory Head (HTH) and Object Affordance Head (OAH). HTH contains an MLP that receives the future parts of the denoised features $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}[N_{\text{p}}\!+\!1\!:\!N_{\text{p}}\!+\!N_{\text{f}}]$ and $\hat{F}^{\scriptstyle{\text{L}}}_{\text{seq}}[N_{\text{p}}\!+\!1\!:\!N_{\text{p}}\!+\!N_{\text{f}}]$, to generate future hand waypoints $\mathcal{H}$ of two hands. As to OAH, we empirically exploit Conditional Variational Autoencoder (C-VAE) [53] to generate possible contact points $\mathcal{O}$ of the next active object in near future. Taking the right hand as an example, the condition is selected as the time-averaged $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}$ and predicted waypoints $H_{t}^{\scriptstyle{\text{R}}}$. Note that we additionally consider denoised future HOI features $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}[N_{\text{p}}\!+\!1:N_{\text{p}}\!+\!N_{\text{f}}]$ (t>0t\!>\!0) besides the features from past observations ($t\!\leq\!0$) for object affordance prediction. This aligns with the intuitive relationship between the contact points and the overall interaction process.
@@ III-C Training
[…]
@@
[…]
@@
[…]
@@
[…]
@@ III-D Inference
[…]
@@ IV-B Separate Evaluation on Hand Trajectory and Object Affordance Prediction
[…]
@@ 1 Motion-Related Gaps and Egomotion Homography
[…]
@@ 4 More Details about Datasets and Diff-IP2D Training Configurations
For training Diff-IP2D, we use AdamW optimizer [63] with a learning rate 22e-44. The total loss function is depicted below. The loss weights are initially set as $\lambda_{\text{VLB}}=1$, $\lambda_{\text{traj}}=1$, $\lambda_{\text{aff}}=0.1$, and $\lambda_{\text{reg}}=0.2$. All the networks in Diff-IP2D are trained for $30$ epochs with a batch size of $8$ on 2 A100 GPUs.